OpenStudy (anonymous):

Usin complete sentences, explain which method u would use 2 solve the following system of equations & why. In ur answer, include the solution 2 1 of the variables & how u found it usin the method u chose. x–5y+2z=0 x+4y–z=12 2x–y+3z=10

OpenStudy (anonymous):

Elimination method. This method is easiest because the first two equations can easily eliminate/ cancel out the X variable. x–5y+2z=0 -x-4y+z=-12 Multiply this equation with -1 in order to make x negative -------------- -9y+3z=-12 New Equation 1 -2x-8y+2z=-24 Multiply this by -2 2x–y+3z=10 ------------ -9y+5z=-14 New Equation 2 Using the two equations that have just been found, use the elimation method once again. +9y-3z=+12 Multiply equation 1 by -1 . -9y+5z=-14 New Equation 2 ------------- 2z = -2 z= -1. Find other variables by substituting it in the equations.